# Quantitative Ultrasound for Periodontal Soft Tissue Characterization

### v1

@@ Abstract
Periodontal (gum) diseases reportedly affect 45.9% of adults 30 years of age or older in the United States. Current diagnostic methods for clinical assessment of periodontal soft tissues are visual examination, palpation which are subjective and qualitative, and bleeding on probing (invasive) which is the late indicator of tissue destruction. Therefore, there is a critical need for research on noninvasive modalities for clinical assessments of periodontal tissues. Quantitative Ultrasound (QUS) analysis has shown promising results in noninvasive characterization of various soft tissues; however, it has not been used in periodontics. This study is among initial investigations into the application of QUS for periodontal tissue characterization in the literature. Here, QUS analysis of oral soft tissues (alveolar mucosa and gingiva) is performed in an in vivo animal study including 10 swine (6 females and 4 males) in which ultrasonic scanning was performed at the first molar of all four oral quadrants, resulting in a total of 40 scans. We investigated first order ultrasonic speckle statistics of oral tissues by using the two-parameter Burr model (power-law $b$ and scale factor $l$) and the two-parameter Nakagami model (shape factor $m$ and scale factor $\alpha$; where $\alpha$ represents the echo intensity). Parametric imaging of these parameters was created using a sliding kernel method sweeping regions of interest with a kernel size of 10 wavelengths chosen from a separate phantom study. Parametric images were superimposed onto the B-mode image to incorporate additional information to its anatomic references and grayscale echogenicity as well as facilitating visual comparisons of the two oral tissue types. Results show that the gingiva and alveolar mucosa are distinct from average Burr and Nakagami parameters. The difference between the two tissue types using model parameters are statistically significant ($p-value$$<$0.0001). Comparing average parameters in swine population, the Burr power-law parameter and Nakagami shape factor are both higher in gingiva than alveolar mucosa while the Burr and Nakagami scale factors are both lower in gingiva. Findings from QUS analyses are in agreement with observation in histology images from the Masson’s Trichrome and hematoxylin–eosin (H&E) staining methods. Both stains show different underlying structure in the two tissue types with gingiva demonstrating denser underlying structure. Linear classifications of these two tissue types using two-dimensional parameter spaces of the Burr and Nakagami models result in a segmentation accuracy of 93.51% and 90.91%, respectively. We propose that QUS holds promising potentials to be employed for the assessment of periodontal soft tissues with the aim of improving diagnostic sciences of periodontology and implant dentistry. QUS could become an objective and quantitative diagnostic tool for the quantification of periodontal soft tissue pathologies and thus, improving dental healthcare.
@@ 1 Introduction
Periodontal (gum) diseases are reported to affect nearly half (45.9%) of the adult populations aged $\geq 30$ years in the United States [1]. These diseases concern various oral soft tissues that support and surround teeth such as marginal (free) gingiva, attached gingiva and, alveolar mucosa, which are illustrated in Figure 1 for a swine model. The most prevalent periodontal diseases affecting these tissues are periodontitis and gingivitis which are a continuum of inflammatory diseases. Periodontitis is initiated by bacteria infection, potentiated by inflammation, resulting in periodontal attachment loss of soft tissues from bone as well as causing actual bone loss. As periodontitis progresses, it could cause tooth loss. It is also related to systemic diseases, such as cardiovascular diseases and diabetes. On the other hand, gingivitis is considered reversible, involving gingival inflammation without clinical signs of bone loss [2]. If oral diseases, affecting both soft and hard tissues, are not addressed at early stages, those could impose immense pain as well as excessive economic burdens on the population, (it is reported to have caused direct and indirect burdens of as high as \$154.06B in the United States in 2018 [3].
@@
Towards this clinical goal, an imaging modality with promising potentials for oral soft tissue characterization is ultrasound (US) imaging. As a non-invasive, non-ionizing, real-time, inexpensive, and well-established modality, US has been employed for imaging and characterization of various biological soft tissues such as liver, thyroid, muscle, etc. at differing depths and image resolutions [7, 8, 9]. In dentistry, US B-mode (brightness-mode) imaging has been employed for lesion detection, measuring gingiva thickness and to delineate the surface of hard tissues (bone/crown) [10, 11, 12, 13]. Moreover, ultrasound-based elasticity estimations of oral soft tissues have been investigated in different studies [14, 15]. Ultrasound imaging offers information beyond B-mode imaging. One important aspect to employ ultrasound imaging is quantitative ultrasound (QUS). In QUS analysis of tissues, the goal is to find quantitative parameters from uncompressed raw ultrasound scan data that can be linked to some measure of underlying structure of tissues, which could offer clinical potentials for tissue characterization [8, 16, 17]. While B-mode images provide information about landmark anatomical structures of tissues, it fails to provide information and contrast of the underlying soft tissue structure. QUS parameters could add more information to B-mode images, represented as a parametric image overlay. Although QUS analysis has been extensively applied to characterize various biological tissues [18, 19, 20, 21, 22], it has never been applied in clinical periodontics for soft tissue characterization and there are only few studies with limited analysis involving QUS for periodontal soft tissues. For example, in the study reported in [23], US B-scan echogenicity in layers of oral soft tissues were compared using a measure of echo levels. Nevertheless, the employed method deviates from standard techniques of analyzing US image echogenicity parameters quantitatively.
@@
One class of QUS analysis in medical imaging of tissues is focused on modeling first order statistics of ultrasound speckle. Speckles are granular (grey) textures observed in B-mode tissue images. Speckles result from the interference of backscattered waves echoed back from various tissue scatterers close to each other during ultrasound pulse-echo imaging. Although speckles have a negative impact on B-mode image quality and are filtered out for US image representation, speckle patterns could incorporate information about underlying tissue structure and thus, could have clinical significance. Modeling ultrasound speckle statistics may result in deriving quantitative parameters that can be correlated to tissue pathology not visible on B-mode images. A number of well-establish distributions for speckle modeling in QUS include Rayleigh [24, 25], Homodyned-K [26, 27, 28], Nakagami [29, 30, 31], and more recently the Burr model [32, 33, 34]. All these distributions have been widely used for QUS-based tissue characterization. However, to the best of our knowledge, characterization of periodontal soft tissues using speckle modeling and these well-established distributions have not been reported in the literature. Here, we aim at characterizing periodontal soft tissue by investigating US speckle statistics using the Burr and Nakagami models in an in vivo animal study on swine oral tissues. Moreover, we present parametric imaging of these QUS parameters as an additional information to that of B-mode images. Additionally, histology images of swine oral tissues were acquired using Masson’s Trichrome and hematoxylin–eosin (H&E) stains with a 20x magnification microscopy imaging to compare tissue structures with QUS analysis. Gingival and alveolar mucosal tissues were compared. QUS-based parametric imaging may have potential to be used as an augmented tool to current imaging modalities in dentistry such as cone beam computed tomography (CBCT) to further aid oral surgeons and to provide diagnostic value to clinical assessment in oral examinations.
@@ 1.1 Oral soft tissues: anatomy
Alveolar mucosa is a membrane that lines the bone. Unlike gingiva, alveolar mucosa is less exposed to abrasive forces and is mainly non-keratinized. It has a higher level of elastic fibers which makes the alveolar mucosa more elastic compared to the gingiva whereas the gingiva contains higher collagen fiber levels cross-linked and possesses some resistance to tensile loads. These elastic fibers tend to make alveolar mucosa return to the resting state when being extended. Moreover, it is distinguished by a higher blood vessel supply, ranging from capillaries to larger ones and appears pinkish compared to the gingiva’s brighter white-pink color [40]. The interstitial fluid within the vasculature provides cushioning when tissue is under large masticatory loads.
@@ 2.1 The Burr model for speckle statistics modeling
The two parameters of $b$ and $l$ have shown potential in characterizing changes in scattering structures of soft tissues such as normal and fibrotic livers [42, 22]. To estimate the Burr parameters from the tissue backscatter within a selected region of interest (ROI), we can fit the echo envelope PDF of the speckle data to equation 1 and derive the underlying parameters. Also, the Burr parameters can be estimated locally from the local statistics of speckle data using a sliding window approach where the ROI is swept by a small kernel and a local estimation map of the Burr parameters is calculated. To do so, we could utilize some relationships between one or multiple statistical moments of the echo amplitude and the Burr parameters to find a system of two equations with two unknown parameters [47]. One statistic to employ is the first moment of the echo amplitude, i.e. mean: denoted as $E[A]$ and reported in equation 2 . The right hand side of this equation is a function of both $b$ and $l$. Other statistics could be the ratio of the square of the first moment of the echo amplitude, $(E[A])^{2}$, to the first moment of the echo intensity, $E[A^{2}]$, as shown in equation 3. The right side of this equation only depends on $b$. Using these two equations, one can obtain local estimations of the Burr parameters.
@@ 2.2 Nakagami Distribution
The probability distribution of the echo amplitude envelope from the two-parameter Nakagami distribution is modeled according to equation 4 where $m$ is the Nakagami shape parameter and $\alpha$ is the Nakagami scale factor. These two parameters are estimated statistically from equation 5 and equation 6. It is noted that the Nakagami $m$ parameter determines the form of the speckle statistics PDF: if $m<1$, it is pre-Rayleigh and a heavy-tailed distribution, if $m=1$, it corresponds to Rayleigh behavior and for $m>1$, it demonstrates a post-Rayleigh distribution [31]. The Nakagami scale factor $\alpha$ represents the total intensity of the backscattered echo within the region under analysis
$f(r)=\frac{2m^{m}r^{2m-1}}{\Gamma(m)\alpha^{m}}e^{-\frac{m}{\alpha}r^{2}}U(r)$ (4)
@@
$\alpha=E[A^{2}]$ (6)
In these equations, $f(r)$ is the Nakagami density function and $U(r)$ is the unit step function.
@@ 3.1 Study design for intraoral US imaging of periodontal soft tissues
For the QUS investigations of periodontal soft tissues in vivo, there are some challenges associated with intraoral US imaging itself which require an intricate study design and customized scanning setup to tackle them [48]. Two most important ones are dealing with the intraoral scanning of inherently small-sized periodontal soft tissues and also the mixed presence of hard and soft tissues. In addition, in intraoral scanning, the buccal (cheek) region imposes a less accommodating condition to foreign objects such as an ultrasound transducer. This limits the capability of freely placing and maneuvering the transducer when scanning the tissue. To address these challenges, we benefited from using a, recently introduced, tooth brush-sized high-frequency US transducer [49]. Its physical design and ability to provide high-resolution images was made for the purpose of intraoral imaging [48] through a collaboration of clinicians and scientists (co-authors H-L.C. and O.D.K) at the University of Michigan (Ann Arbor, MI) and the Mindray Innovation Center (Mindray Inc., San Jose, CA). This US transducer is shown in Figure 2 (a) and (b), with its in situ placement in Figure 2 (c). In order to acquire high-quality gingival and mucosal scans, a standoff gel pad was placed onto the aperture surface to shift the tissues of interest to the focal region of the transducer, as shown in Figure 2 (d). Having in mind the anatomical structure of oral tissues shown in Figure 2 (e), a sample of US scan of periodontal soft tissues in a swine model is presented in Figure 2 (f). Regions of the gingiva, mucosa, epithelium, muscle as well as crown and bone are annotated to help distinguish regions of interest for QUS analysis. It is noted that in the B-scan presented in (f), the transducer is located on the right edge of the image. The B-scan sample shows that the epithelium appears as a hyper-echoic (bright) region, indicating a higher US backscatter intensity from this region. This higher scattering phenomenon may result from contributions of dense keratinized tissue (higher impedance). Also, the normal orientation of the stratified epithelium layer with respect to the transmitted pulse from the US probe may produce the strongest reflection compared to other angles.
@@
Swine models were selected for this study due to the resemblance of human and swine oral soft tissues from histology and morphology aspects [50]. Animals (N=10) were obtained from Sinclair Bio Resources (Auxvasse, MO, USA) under a study protocol approved by the Institutional Animal Care and Use Committee (PRO00010333). Four males and six females were investigated. Intraoral US scanning was performed at the mid-facial location of all four first molars (M1), which we refer to by combinations of L (left) / R (right) and mandible (MAND) / maxilla (MAX). It is noted that a small notch was manually created on all teeth as a landmark visible on ultrasound images.

### v2

@@ Abstract
Objective
Periodontal diseases affect 45.9% of adults of age $\geq 30$ years in the United States. Current clinical diagnostic methods are invasive, subjective and qualitative. Quantitative Ultrasound (QUS) has shown promising results in noninvasive characterization of various soft tissues; however, it has not been used in clinical periodontics.
Methods
Here, we investigated QUS analysis of two periodontal soft tissues (alveolar mucosa and gingiva) in vivo. Study cohort included 10 swine scanned at four oral quadrants, resulting in 40 scans. The two-parameter Burr and Nakagami were employed for QUS-based speckle modeling. Parametric imaging of these parameters was also created using an optimal estimation window size (WS) in a separate phantom study.
Results
Phantom results suggested a WS of 10 wavelengths as the reasonable estimation kernel. The Burr power-law parameter and Nakagami shape factor were higher in gingiva than alveolar mucosa while the Burr and Nakagami scale factors were lower in gingiva. The difference between the two tissue types was statistically significant (p$<$0.0001). Linear classifications of these two tissue types using a 2D parameter space of the Burr and Nakagami models resulted in a segmentation accuracy of 93.51% and 90.91%, respectively. Findings from histology-stained images showed that gingiva and alveolar mucosa had distinct underlying structure with gingiva showing a denser stain.
Conclusion
QUS results suggested that the gingiva and alveolar mucosa were differentiable using the Burr and Nakagami parameters. We propose that QUS holds promising potential for characterization of periodontal soft tissues. QUS could become an objective and quantitative diagnostic tool for periodontology and implant dentistry to improve dental healthcare.
@@ 1 Introduction
Periodontal (gum) diseases are reported to affect nearly half (45.9%) of the adult population aged $\geq 30$ years in the United States [1]. These diseases concern various oral soft tissues that support and surround teeth such as marginal (free) gingiva, attached gingiva and, alveolar mucosa, which are illustrated in Figure 1 for a swine model. The most prevalent periodontal diseases affecting these tissues are periodontitis and gingivitis which are a continuum of inflammatory diseases. Periodontitis is initiated by bacteria infection, potentiated by inflammation, resulting in periodontal attachment loss of soft tissues from bone as well as causing actual bone loss. As periodontitis progresses, it could cause tooth loss. It is also related to systemic diseases, such as cardiovascular diseases and diabetes. On the other hand, gingivitis is considered reversible, involving gingival inflammation without clinical signs of bone loss [2]. If oral diseases, affecting both soft and hard tissues, are not addressed at early stages, those could impose immense pain as well as excessive economic burdens on the population, (it is reported to have caused direct and indirect burdens of as high as \$154.06B in the United States in 2018 [3].
@@
Towards this clinical goal, an imaging modality with promising potentials for oral soft tissue characterization is ultrasound (US) imaging. As a non-invasive, non-ionizing, real-time, inexpensive, and well-established modality, US has been employed for imaging and characterization of various biological soft tissues such as liver, thyroid, muscle, etc. at differing depths and image resolutions [7, 8, 9]. In dentistry, US B-mode (brightness-mode) imaging has been employed for lesion detection, measuring gingiva thickness and to delineate the surface of hard tissues (bone/crown) [10, 11, 12, 13]. Moreover, ultrasound-based elasticity estimations of oral soft tissues have been investigated in different studies [14, 15]. Ultrasound imaging offers information beyond B-mode imaging. One important aspect to employ ultrasound imaging is quantitative ultrasound (QUS). In QUS analysis of tissues, the goal is to find quantitative parameters from uncompressed raw ultrasound scan data that can be linked to some measure of underlying structure of tissues, which could offer clinical potentials for tissue characterization [8, 16, 17]. While B-mode images provide information about landmark anatomical structures of tissues, it fails to provide information and contrast of the underlying soft tissue structure. QUS parameters could add more information to B-mode images, represented as a parametric image overlay. Although QUS analysis has been extensively applied to characterize various biological tissues [18, 19, 20, 21, 22], it has never been applied in clinical periodontics for soft tissue characterization and there are only few studies with limited analysis involving QUS for periodontal soft tissues. For example, in the study reported in [23], US B-scan echogenicity in layers of oral soft tissues were compared using a measure of echo levels. Nevertheless, the employed method deviates from standard techniques of analyzing US image echogenicity parameters quantitatively. Potential application of QUS in periodontology includes tissue characterization, i.e., gingiva versus a. mucosa, tracking inflammation progress, automatic lesion detection, tissue healing quantification, etc. QUS could potentially complement BOP and other diagnostic tools for an improved and more accurate characterization of periodontal tissues.
@@
One class of QUS analysis in medical imaging of tissues is focused on modeling first order statistics of ultrasound speckle. Speckles are granular (grey) textures observed in B-mode tissue images [24]. Speckles result from the interference of backscattered waves echoed back from various tissue scatterers close to each other during ultrasound pulse-echo imaging. Although speckles have a negative impact on B-mode image quality and are filtered out for US image representation, speckle patterns could incorporate information about underlying tissue structure and thus, could have clinical significance. Modeling ultrasound speckle statistics may result in deriving quantitative parameters that can be correlated to tissue pathology not visible on B-mode images. A number of well-establish distributions for speckle modeling in QUS include Rayleigh [25, 26], Homodyned-K [27, 28, 29], Nakagami [30, 31, 32], and more recently the Burr model [33, 34, 35]. All these distributions have been widely used for QUS-based tissue characterization. However, to the best of our knowledge, characterization of periodontal soft tissues using speckle modeling and these well-established distributions have not been reported in the literature. Here, we aim at characterizing periodontal soft tissue by investigating US speckle statistics using the Burr and Nakagami models in an in vivo animal study on swine oral tissues. Moreover, we present parametric imaging of these QUS parameters as an additional information to that of B-mode images. Additionally, histology images of swine oral tissues were acquired using Masson’s Trichrome and hematoxylin–eosin (H&E) stains with a 20x magnification microscopy imaging to compare tissue structures with QUS analysis. Gingival and alveolar mucosal tissues were compared. QUS-based parametric imaging may have potential to be used as an augmented tool to current imaging modalities in dentistry such as cone beam computed tomography (CBCT) to further aid oral surgeons and to provide diagnostic value to clinical assessment in oral examinations.
@@ 1.1 Oral Soft Tissues: Anatomy
Alveolar mucosa is a membrane that lines the bone. Unlike gingiva, alveolar mucosa is less exposed to abrasive forces and is mainly non-keratinized. It has a higher level of elastic fibers which makes the alveolar mucosa more elastic compared to the gingiva. The gingiva contains higher collagen fiber levels cross-linked and possesses some resistance to tensile loads. These elastic fibers tend to make alveolar mucosa return to the resting state when being extended. Moreover, it is distinguished by a higher blood vessel supply, ranging from capillaries to larger ones and appears pinkish compared to the gingiva’s brighter white-pink color [41]. The interstitial fluid within the vasculature provides cushioning when tissue is under large masticatory loads.
@@ 2.1 The Burr Model for Speckle Statistics Modeling
Burr b is associated with the number density of scatterers and increases with that while Burr l is related to the echo amplitude and is elevated in tissues with higher echogenicity. These parameters provide additional information about the tissue’s underlying structure compared to the grey-scale B-scan alone. Burr b and l have shown potential in characterizing liver tissues in fibrotic vs. normal conditions [22, 43]. To estimate the Burr parameters from the tissue backscatter within a selected region of interest (ROI), we can fit the echo envelope PDF of the speckle data to equation  1 and derive the underlying parameters. Also, the Burr parameters can be estimated locally from the local statistics of speckle data using a sliding window approach where the ROI is swept by a small kernel and a local estimation map of the Burr parameters is calculated. To do so, we could utilize relationships between one or multiple statistical moments of the echo amplitude and the Burr parameters to find a system of two equations with two unknown parameters [48]. One statistic to employ is the first moment of the echo amplitude, i.e. mean: denoted as $E[A]$ and reported in equation 2. Other statistics could be the ratio of the square of the first moment of the echo amplitude, $(E[A])^{2}$, to the first moment of the echo intensity, $E[A^{2}]$, as shown in equation 3. While equation 2 shows that the first moment depends on both b and l, per equation  3 the moments’ ratio depends only on one single parameter, (b). To estimate these two parameters, first b is obtained from Equation 3 and then, it is plugged into equation 2 to estimate l. The Burr scale factor, l, in equation  2 is constrained by the finite value of b (measure of the scatterer number density, from equation  3) and by the value of the first moment of the amplitude (mean), E[A]. Thus, l does not diverge in practice due to the finite nature of both the backscatter amplitude and the b parameter. Using these two equations, one can obtain local estimations of the Burr parameters.
@@ 2.2 Nakagami Distribution
The probability distribution of the echo amplitude envelope from the two-parameter Nakagami distribution is modeled according to equation 4 where $m$ is the Nakagami shape parameter and $\Omega$ is the Nakagami scale factor. These two parameters are estimated statistically from equation 5 and equation 6. It is noted that the Nakagami $m$ parameter determines the form of the speckle statistics PDF: if $m<1$, it is pre-Rayleigh and a heavy-tailed distribution, if $m=1$, it corresponds to Rayleigh behavior and for $m>1$, it demonstrates a post-Rayleigh distribution [32]. The Nakagami scale factor $\Omega$ represents the total intensity of the backscattered echo within the region under analysis.
$f(r)=\frac{2m^{m}r^{2m-1}}{\Gamma(m)\Omega^{m}}e^{-\frac{m}{\Omega}r^{2}}U(r)$ (4)
@@
$\Omega=E[A^{2}]$ (6)
In these equations, $f(r)$ is the Nakagami density function and $U(r)$ is the unit step function. The Nakagami shape factor m determines the form of the probability distribution function of speckle data and is related to the scatterer concentration while the Nakagami scale factor $\Omega$ quantifies the backscatter intensity [32]. These parameters have been used in various studies for tissue characterization. For example, in an animal model study reported in [31], it was shown that the Nakagami shape factor increased with increase in the liver fibrosis score in rats. This was postulated to result from the increase in the concentration (density) of local scatterers and also the appearance of periodic structures or clustering of scatterers in tissues, leading to a post-Rayleigh statistical distribution and higher m.
@@ 3.1 Study Design for Intraoral US Imaging of Periodontal Soft Tissues
For the QUS investigations of periodontal soft tissues in vivo, there are some challenges associated with intraoral US imaging itself which require an intricate study design and customized scanning setup to tackle them [49]. Two most important ones are dealing with the intraoral scanning of inherently small-sized periodontal soft tissues and also the mixed presence of hard and soft tissues. In addition, in intraoral scanning, the buccal (cheek) region imposes a less accommodating condition to foreign objects such as an ultrasound transducer. This limits the capability of freely placing and maneuvering the transducer when scanning the tissue. To address these challenges, we benefited from using a, recently introduced, toothbrush-sized high-frequency US transducer [50]. Its physical design and ability to provide high-resolution images was made for the purpose of intraoral imaging [49] through a collaboration of clinicians and scientists (co-authors H-L.C. and O.D.K) at the University of Michigan (Ann Arbor, MI) and the Mindray Innovation Center (Mindray Inc., San Jose, CA). This US transducer is shown in Figure 2 (a) and (b), with its in situ placement in Figure 2 (c). The center frequency of the transducer is 18 MHz, the imaging depth and the lateral field of view commonly used are approximately 15 mm and 13 mm, respectively and the transducer elevational focus is at a depth of 8 mm. The useable bandwidth of this array is 8 to 30 MHz. The system was operated in a single transmit and receive frequency mode for the QUS data acquisition in this study. The pre-clinical images (not IQ data) were obtained in the standard second harmonic imaging mode (12 MHz transmit, 24 MHz receive) with additional spatial compounding to reduce clutter and for obtaining high-resolution ultrasound images. In order to acquire high-quality gingival and mucosal scans, a standoff gel pad was placed onto the aperture surface to shift the tissues of interest to the focal region of the transducer, as shown in Figure 2 (d). Having in mind the anatomical structure of oral tissues shown in Figure 2 (e), a sample of US scan of periodontal soft tissues in a swine model is presented in Figure 2 (f). Regions of the gingiva, mucosa, epithelium, muscle as well as crown and bone are annotated to help distinguish regions of interest for QUS analysis. It is noted that in the B-scan presented in (f), the transducer is located on the right edge of the image. The B-scan sample shows that the epithelium appears as a hyper-echoic (bright) region, indicating a higher US backscatter intensity from this region. This higher scattering phenomenon may result from contributions of dense keratinized tissue (higher impedance). Also, the normal orientation of the stratified epithelium layer with respect to the transmitted pulse from the US probe may produce the strongest reflection compared to other angles.
@@
Swine models were selected for this study due to the resemblance of human and swine oral soft tissues from histology and morphology aspects [51]. Animals (N=10) were obtained from Sinclair Bio Resources (Auxvasse, MO, USA) under a study protocol approved by the Institutional Animal Care and Use Committee at the University of Michigan (PRO00010333). Four males and six females were investigated. Intraoral US scanning was performed at the mid-facial location of all four first molars (M1), which we refer to by combinations of L (left) / R (right) and mandible (MAND) / maxilla (MAX). It is noted that a small notch was manually created on all teeth as a landmark visible on ultrasound images.